OpenStudy (anonymous):

Solve 3x2 = -12x - 15. Please helppppppppp!

OpenStudy (anonymous):

A.) x = -4 ± 2i B.) x = -4 ± i C.) x = -2 ± 2i D.) x = -2 ± i

OpenStudy (doc.brown):

\[3x^2 = -12x-15\]Put it in the form\[ax^2+bx+c=0\]

OpenStudy (doc.brown):

Solve for x using the quadratic formula\[x=\frac{b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

When plugging in the first part would i chnge the sign of -12 or?

OpenStudy (anonymous):

For example would it be 3x2+12+15=0?

OpenStudy (doc.brown):

Change the sign... I guess you could think of it like that. What I would do is add 12x to both sides of the equation, to keep it equal.\[3x^2\color{red}{+12x}= -12x\color{red}{+12x}-15\]\[3x^2+12x= 15\]then add 15 to both sides\[3x^2+12x\color{red}{+15}= -15\color{red}{+15}\]\[3x^2+12x+15=0\]

OpenStudy (doc.brown):

sorry, in that first step i meant =-15

OpenStudy (anonymous):

So would i use the quadratic formula now?

OpenStudy (doc.brown):

yes

OpenStudy (anonymous):

so it would be